OpenStudy (anonymous):

what is an equation of the line in point-slope form that passes through the given point and has the given slope? point: (9,3); slope: 4/9 a.) y-9 = 4/9(x+3) b.) y-9 = 4/9(x-3) c.) y+3 = 4/9(x-9) d.) y-3 = 4/9(x-9)

OpenStudy (anonymous):

Tried this but cant seem to get it, would appreciate some assistance.

OpenStudy (jdoe0001):

http://zonalandeducation.com/mmts/functionInstitute/linearFunctions/psPic.gif plug the (x,y) values there, along with the slope

OpenStudy (anonymous):

So it would look like y - 9(1) = 4/9(x-3(1))?

OpenStudy (jdoe0001):

well, nevermind the "1" in the point-slope graphic there, it just stands for $$ y_2-y_1=m(x_2-x_1) $$

OpenStudy (anonymous):

So the answer would be B correct?

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

sorry was away, no, recheck the values :)